a woman's $57,000 is to be divided so that her husband receives twice as much as her son. how much will each receive? set up and equation and then solve.
Answer by robertb(4012)

s = portion of the son.
57000 - s = portion of the father. Then
2s = 5700 - s,
3s = 57000,
s = 19000, portion of the son,
57000-19000 = 38000, portion of the father.